# Quantum metrology with a continuous-variable system

### v1

@@ II.1 Quadrature operators and Wigner function
The quantum description of a single-mode continuous variable system is equivalent to that of a quantum harmonic oscillator. The position and momentum operators $\hat{x}$ and $\hat{p}$ satisfy the canonical commutation relations $[\hat{x},\hat{p}]=i$, where henceforth we will set $\hbar=1$. We introduce bosonic creation and annihilation operators $\hat{a}$ and $\hat{a}^{\dagger}$, satisfying $[\hat{a},\hat{a}^{\dagger}]=1$, such that
@@
where $|x\rangle$ are the eigenstates of the position operator $\hat{x}$. Normalization implies $\int\text{d}x\,\text{d}p\,W(x,p)=1$. The information contained in the Wigner function is equivalent to that of the full quantum state $\hat{\rho}$ and therefore provides a complete description of the CV system’s state. We will often switch between the phase space coordinates $(x,p)\in\mathbb{R}^{2}$ and $\alpha\in\mathbb{C}$ using the mapping $\alpha=\frac{x+ip}{\sqrt{2}}$. The Wigner function can be written as the expectation value of a displaced parity measurement [61], namely
@@
where the factor 2 comes from the Jacobian of the coordinate transformation, such that $\int\text{d}^{2}\alpha\,W(\alpha)=1$.
@@ II.2 Quantum states of interest
where $\hat{S}(\xi)$ is the squeezing operator
@@
with $\hat{\rho}_{T}$ the thermal state with average number of thermal excitations $n_{T}\equiv\text{Tr}[\hat{\rho_{T}}\hat{a}^{\dagger}\hat{a}]$. Here, note that the total average number of excitation in the state is $\overline{n}\equiv\text{Tr}[\hat{\rho}(\alpha,\xi,n_{T})\hat{a}^{\dagger}\hat{a}]=n_{T}\cosh(2r)+|\alpha|^{2}+\sinh^{2}r$, meaning that for $|\alpha|^{2}=0$ we have $\cosh(2r)=(1+2\overline{n})/(1+2n_{T})$ and $\sinh^{2}r=(\overline{n}-n_{T})/(1+2n_{T})$. Gaussianity implies that these states can be fully described by first and second moments of quadrature operators. For this reason, a Gaussian state is fully determined by its first-moment vector $\langle\mathbf{\hat{r}}\rangle_{\hat{\rho}}$ and its covariance matrix $\Gamma[\hat{\rho},\mathbf{\hat{r}}]$, with $\mathbf{\hat{r}}=(\hat{x},\hat{p})^{T}$, $\langle\hat{M}\rangle_{\hat{\rho}}=\,\mathrm{Tr}\{\hat{M}\hat{\rho}\}$, and
@@ II.3 Gaussian evolutions
Of patricular interest for Gaussian states are Gaussian evolutions, since they maintain the Gaussianity of the state at all times. For a detailed discussion we refer to the extensive literature on the topic [53, 47, 54, 55, 56, 57, 58, 59]. An example of a Gaussian evolution is a unitary evolution generated by a quadratic Hamiltonian
$\hat{U}(\theta)=\exp\left(-i\frac{1}{2}\hat{\bm{r}}^{T}H\hat{\bm{r}}\theta\right).$ (18)
@@
$\mathcal{S}(\theta)=\exp(\Omega H\theta)$ (20)
describes a symplectic transformation and
$\Omega=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$ (21)
is the symplectic form.
@@ III Quantum metrology
Figure 2: Schematic illustration of a metrological protocol. A parameter $\theta$ is imprinted on an initial quantum state $\hat{\rho}$ through the unitary evolution $\hat{U}(\theta)$ generated by the operator $\hat{G}$. This results in the state $\hat{\rho}(\theta)$, on which a measurement $\hat{M}$ is performed to estimate the unknown $\theta$.
Consider a parameter $\theta$ that is imprinted by a unitary evolution $\hat{U}(\theta)=\exp[-i\hat{G}\theta]$, generated by $\hat{G}$. Information about this parameter is imprinted into the system with initial state $\hat{\rho}$ through the evolution $\hat{\rho}(\theta)=\hat{U}(\theta)\hat{\rho}\hat{U}(\theta)^{\dagger}$. The value of $\theta$ can now be estimated from measurements of the average value of an observable $\hat{M}$, see Fig. 2.
@@ III.1 Quantum Fisher information and Cramér-Rao bound
The Cramér-Rao (CR) bound defines a fundamental limit for the standard deviation of an unbiased estimator $\theta_{\mathrm{est}}$ for $\theta$:
$(\Delta\theta_{\mathrm{est}})^{2}\geq\frac{1}{\mu F[\hat{\rho}(\theta),\hat{G},\{\hat{E}_{m}\}]}\;,$ (23)
@@
$F[\hat{\rho}(\theta),\hat{G},\{\hat{E}_{m}\}]=\sum_{m}\frac{1}{P(m|\theta)}\left(\frac{\partial P(m|\theta)}{\partial\theta}\right)^{2}\;$ (24)
is the Fisher information. This quantity describes the information that can be retrieved about the parameter $\theta$ from measurement results $m$ that are distributed according to the probability distribution $P(m|\theta)$, conditioned on $\theta$. A measurement on a quantum system is described by a positive operator-valued measure (POVM), i.e. a set $\{\hat{E}_{m}\}_{m}$ of positive semidefinite operators $\hat{E}_{m}\geq 0$ such that $\sum_{m}\hat{E}_{m}=\hat{\mathbb{I}}$. The probability of obtaining the result $m$ given a phase $\theta$ is then given, according to quantum theory, by $P(m|\theta)=\,\mathrm{Tr}[\hat{\rho}(\theta)\hat{E}_{m}]$, where $\hat{\rho}(\theta)$ is the quantum state of the system. It can be shown that a maximum likelihood analysis of the measurement data saturates the CR bound when a large enough sample has been recorded.
Since the Fisher information is a function of the measurement setting, the achievable precision still depends crucially on the chosen POVM. A POVM of particular interest is the one that minimizes the estimation error and gives rise to the quantum Fisher information (QFI):
$F_{\mathrm{Q}}[\hat{\rho},\hat{G}]=\max_{\{\hat{E}_{m}\}}F[\hat{\rho}(\theta),\hat{G},\{\hat{E}_{m}\}]\;,$ (25)
For unitary evolutions, the QFI is independent of $\theta$ and only depends on the initial state $\hat{\rho}$ and the generator $\hat{G}$. In full generality, this can be obtained as
$F_{\mathrm{Q}}[\hat{\rho},\hat{G}]=2\sum_{k,l}\frac{(\lambda_{k}-\lambda_{l})^{2}}{(\lambda_{k}+\lambda_{l})}|\langle k|\hat{G}|l\rangle|^{2}\;,$ (26)
where $\lambda_{k}$ and $|k\rangle$ are the eigenvalues and eigenvectors of $\hat{\rho}$, respectively, and the summation goes over all $k,l$ such that $\lambda_{k}+\lambda_{l}>0$. In many relevant situations, however, this expression can be significantly simplified.
@@
In addition, an analytical expression for the QFI can also be obtained for arbitrary Gaussian states under unitary evolutions. For a single-mode Gaussian state $\hat{\rho}(\theta)$ under unitary evolution, the QFI is given in terms of their first-moment vector $\langle\mathbf{\hat{r}}\rangle_{\hat{\rho}(\theta)}$ and covariance matrix $\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}]$ as [33, 64, 65, 58]
$F_{Q}[\hat{\rho}(\theta)]=\frac{1}{2}\frac{\,\mathrm{Tr}\{(\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\frac{\partial}{\partial\theta}\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}])^{2})\}}{1+\mathcal{P}(\hat{\rho}(\theta))^{2}}+\left(\frac{\partial\langle\mathbf{\hat{r}}\rangle_{\hat{\rho}(\theta)}}{\partial\theta}\right)^{T}\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\left(\frac{\partial\langle\mathbf{\hat{r}}\rangle_{\hat{\rho}(\theta)}}{\partial\theta}\right),$ (29)
@@
is the purity. Note that for non-unitary evolution, an additional term appears that depends on the derivative of the purity [33, 58]. For an evolution generated by a quadratic Hamiltonian of the form (18), we obtain
$F_{Q}[\hat{\rho},\frac{1}{2}\hat{\bm{r}}^{T}H\hat{\bm{r}}]$ $=\frac{\,\mathrm{Tr}\{(\Gamma[\hat{\rho},\mathbf{\hat{r}}]H)^{2}\}+\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]\,\mathrm{Tr}\{(\Omega H)^{2}\}}{\frac{1}{4}+\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}+\frac{\langle\hat{\bm{r}}\rangle_{\hat{\rho}}^{T}H\Gamma[\hat{\rho},\mathbf{\hat{r}}]H\langle\hat{\bm{r}}\rangle_{\hat{\rho}}}{\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}.$ (31)
A derivation of this result is given in App. F.
The QFI is a convex function of the quantum state, i.e. for any probability distribution $p_{k}$ and family of quantum states $\hat{\rho}_{k}$, we have
@@ III.2 Method of moments
The QFI describes the maximal metrological sensitivity that can be extracted from the quantum state $\hat{\rho}$ using an optimal estimation strategy. A practical approach to parameter estimation based on the average value of a single observable $\hat{M}$ is given by the method of moments. This method provides a simple estimator for $\theta$ by comparing the sample average of $\mu$ repeated measurements of $\hat{M}$ with a previously established relation $\langle\hat{M}\rangle_{\hat{\rho}(\theta)}$, i.e. a “calibration curve”. Once this relation is known, implementation of the method of moments is straightforward: The estimator is simply given by the parameter value that yields equality between the sample average and the calibration curve.
@@
Dichotomic measurements. For dichotomic measurements the method of moments equals the classical Fisher information. Consider an observable $\hat{M}=m_{1}\hat{\Pi}_{1}+m_{2}\hat{\Pi}_{2}$ with two eigenvalues $m_{i}$ and corresponding projectors $\hat{\Pi}_{i}$, such that $\hat{\Pi}_{1}+\hat{\Pi}_{2}=\hat{\mathbb{I}}$. The conditional probability is given by $P(m_{1}|\theta)=\,\mathrm{Tr}\{\hat{\rho}(\theta)\hat{\Pi}_{1}\}=1-\,\mathrm{Tr}\{\hat{\rho}(\theta)\hat{\Pi}_{2}\}=1-P(m_{2}|\theta)$. The variance of $\hat{M}$ reads
@@
$F[\hat{\rho}(\theta),\hat{G},\{\hat{E}_{m}\}]$ $=\frac{1}{P(m_{1}|\theta)}\left(\frac{\partial P(m_{1}|\theta)}{\partial\theta}\right)^{2}+\frac{1}{P(m_{2}|\theta)}\left(\frac{\partial P(m_{2}|\theta)}{\partial\theta}\right)^{2}$

### v2

@@ II.1 Quadrature operators and Wigner function
The quantum description of a single-mode continuous variable system is equivalent to that of a quantum harmonic oscillator. The position and momentum operators $\hat{x}$ and $\hat{p}$ satisfy the canonical commutation relations $\hat{x},\hat{p}]=i$, where henceforth we will set $\hbar=1$. We introduce bosonic creation and annihilation operators $\hat{a}$ and $\hat{a}^{\dagger}$, satisfying $\hat{a},\hat{a}^{\dagger}]=1$, such that
@@
where $|x\rangle$ are the eigenstates of the position operator $\hat{x}$. Normalization implies $\int\text{d}x\,\text{d}p\,W(x,p)=1$. The information contained in the Wigner function is equivalent to that of the full quantum state $\hat{\rho}$ and therefore provides a complete description of the CV system’s state. We will often switch between the phase space coordinates $(x,p)\in\mathbb{R}^{2}$ and $\alpha\in\mathbb{C}$ using the mapping $\alpha=\frac{x+ip}{\sqrt{2}}$. The Wigner function can be written as the expectation value of a displaced parity measurement [61, 62], namely
@@
where the factor 2 comes from the Jacobian of the coordinate transformation, such that $\int\text{d}^{2}\alpha\,W(\alpha)=1$. Equation (2) provides a prescription to experimentally measure the Wigner function when parity measurements are available, such as in cavity QED and trapped ions [63]. In other cases, however, parity measurements can be challenging and one needs to rely on other observables. For example, typical measurement performed on propagating optical fields are homodyne measurements, from which the Wigner function can be reconstructed via tomographic approaches [64].
@@ II.2 Quantum states of interest
where $\hat{S}(\xi)$ is the squeezing operator [65, 66, 67, 68]
@@
with $\hat{\rho}_{T}$ the thermal state with average number of thermal excitations $n_{T}\equiv\text{Tr}[\hat{\rho_{T}}\hat{a}^{\dagger}\hat{a}$]. Here, note that the total average number of excitation in the state is $\overline{n}\equiv\text{Tr}[\hat{\rho}(\alpha,\xi,n_{T})\hat{a}^{\dagger}\hat{a}]=n_{T}\cosh(2r)+|\alpha|^{2}+\sinh^{2}r$, meaning that for $|\alpha|^{2}=0$ we have $\cosh(2r)=(1+2\overline{n})/(1+2n_{T})$ and $\sinh^{2}r=(\overline{n}-n_{T})/(1+2n_{T})$. Gaussianity implies that these states can be fully described by first and second moments of quadrature operators. For this reason, a Gaussian state is fully determined by its first-moment vector $\langle\mathbf{\hat{r}}\rangle_{\hat{\rho}}$ and its covariance matrix $\Gamma[\hat{\rho},\mathbf{\hat{r}}$], with $\mathbf{\hat{r}}=(\hat{x},\hat{p})^{T}$, $\langle\hat{M}\rangle_{\hat{\rho}}=\,\mathrm{Tr}\left[\hat{M}\hat{\rho}\right$], and
@@
A necessary and sufficient condition for a matrix $\Gamma$ to describe a Gaussian quantum state is that it satisfies the uncertainty relation $\Gamma+\frac{i}{2}\Omega\geq 0$, where $\Omega_{ij}=[\hat{r}_{i},\hat{r}_{j}]$ is the symplectic form with entries
$\Omega=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$. (16)
For a single-mode system, the uncertainty relation is equivalent to the two conditions $\Gamma\geq 0$ and $\det\Gamma\geq 1/4$ [58], with pure Gaussian states satisfying $\det\Gamma=1/4$.
@@ II.3 Gaussian evolutions
Unitary evolutions. Of particular interest for Gaussian states are Gaussian evolutions, since they maintain the Gaussianity of the state at all times. For a detailed discussion we refer to the extensive literature on the topic [53, 47, 54, 55, 56, 57, 58, 59]. An example of a Gaussian evolution is a unitary evolution generated by a quadratic Hamiltonian
$\hat{U}(\theta)=\exp\left[-i\frac{1}{2}\hat{\bm{r}}^{T}H\hat{\bm{r}}\theta\right].$ (19)
@@
$\mathcal{S}(\theta)=\exp[\Omega H\theta]$ (21)
describes a symplectic transformation with the defining property $\mathcal{S}(\theta)\Omega\mathcal{S}(\theta)^{T}=\Omega$.
@@
Gaussian channels. When the system of interest interacts with an inaccessible environment, its effective evolution is no longer unitary but instead governed by a completely positive trace-preserving map, commonly referred to as a quantum channel [71]. Gaussian channels preserve the Gaussianity of the initial state and are described by the evolution [58]
$\langle\hat{\bm{r}}\rangle_{\hat{\rho}(\theta)}$ $=\mathcal{X}(\theta)\langle\hat{\bm{r}}\rangle_{\hat{\rho}}+\bm{d}(\theta)$
$\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}$] $=\mathcal{X}(\theta)\Gamma[\hat{\rho},\mathbf{\hat{r}}]\mathcal{X}(\theta)^{T}+\mathcal{Y}(\theta),$ (23)
where $\mathcal{X}(\theta)$ and $\mathcal{Y}(\theta)$ satisfy
$\mathcal{Y}(\theta)+\frac{i}{2}\Omega\geq\frac{i}{2}\mathcal{X}(\theta)\Omega\mathcal{X}(\theta)^{T}.$ (24)
These channels will play a crucial role when analyzing the metrological performances of CV systems subject to common decoherence channels, such as losses or heating.
@@ III Quantum metrology
Figure 2: Schematic illustration of a metrological protocol. A parameter $\theta$ is imprinted on an initial quantum state $\hat{\rho}$ through the evolution described by the quantum channel $\Lambda_{\theta}$. This results in the state $\hat{\rho}(\theta)$, on which a POVM measurement $\{E_{m}\}_{m}$ is performed to estimate the unknown $\theta$.
@@ III.1 Quantum Fisher information and Cramér-Rao bound
Classical bound. We consider the task of estimating the value of an unknown parameter that determines the state $\hat{\rho}(\theta)$ of a quantum system, see Fig. 2. To infer the value of $\theta$, we perform repeated measurements on the system and construct an estimator function $\theta_{\mathrm{est}}$ that maps the measurement outcomes onto a possible value for $\theta$. The Cramér-Rao (CR) bound defines a fundamental limit for the variance of an unbiased estimator:
$(\Delta\theta_{\mathrm{est}})^{2}\geq\frac{1}{\mu F[\hat{\rho}(\theta),\{\hat{E}_{m}\}]}\;,$ (25)
@@
$F[\hat{\rho}(\theta),\{\hat{E}_{m}\}]=\sum_{m}\frac{1}{P(m|\theta)}\left(\frac{\partial P(m|\theta)}{\partial\theta}\right)^{2}\;$ (26)
is the Fisher information. This quantity describes the information that can be retrieved about the parameter $\theta$ from measurement results $m$ that are distributed according to the probability distribution $P(m|\theta)$, conditioned on $\theta$. A measurement on a quantum system is described by a positive operator-valued measure (POVM), i.e. a set $\{\hat{E}_{m}\}_{m}$ of positive semidefinite operators $\hat{E}_{m}\geq 0$ such that $\sum_{m}\hat{E}_{m}=\hat{\mathbb{I}}$. The probability of obtaining the result $m$ given a phase $\theta$ is then given, according to quantum theory, by $P(m|\theta)=\,\mathrm{Tr}[\hat{\rho}(\theta)\hat{E}_{m}]$, where $\hat{\rho}(\theta)$ is the quantum state of the system. It can be shown that a maximum likelihood analysis of the measurement data saturates the CR bound when a large enough sample has been recorded [72, 73].
Quantum bound. Since the Fisher information is a function of the measurement setting, the achievable precision still depends crucially on the chosen POVM. A POVM of particular interest is the one that minimizes the estimation error and gives rise to the quantum Fisher information (QFI):
$F_{\mathrm{Q}}[\hat{\rho}(\theta)]=\max_{\{\hat{E}_{m}\}}F[\hat{\rho}(\theta),\{\hat{E}_{m}\}]\;,$ (27)
The QFI can be expressed as
$F_{\mathrm{Q}}[\hat{\rho}(\theta)]=\sum_{k,l}\frac{2}{\lambda_{k}+\lambda_{l}}|\langle k|\frac{\partial\hat{\rho}(\theta)}{\partial\theta}|l\rangle|^{2}\;,$ (28)
where $\lambda_{k}$ and $|k\rangle$ are the eigenvalues and eigenvectors of $\hat{\rho}(\theta)$, respectively, and the summation includes all $k,l$ such that $\lambda_{k}+\lambda_{l}>0$. The optimal POVM that achieves the maximum in Eq. (27) can be obtained as the projectors onto the eigenvalues of the symmetric logarithmic derivative operator, i.e., if $\hat{L}=\sum_{m}l_{m}\hat{E}_{m}$, where $\hat{L}$ is defined as the operator satisfying the relation [74, 1]
$\frac{\partial\hat{\rho}(\theta)}{\partial\theta}=\frac{1}{2}\left(\hat{L}(\theta)\hat{\rho}(\theta)+\hat{\rho}(\theta)\hat{L}(\theta)\right)\;.$ (29)
The QFI can be expressed using $\hat{L}(\theta)$ as $F_{\mathrm{Q}}[\hat{\rho}(\theta)]=\,\mathrm{Tr}{\left[\hat{\rho}(\theta)\hat{L}(\theta)^{2}\right]}$.
Unitary evolution. A relevant special scenario is that of a unitarily imprinted parameter. A unitary evolution $\hat{U}(\theta)=\exp[-i\hat{G}\theta]$, generated by $\hat{G}$ transforms the initial state $\hat{\rho}$ into the parameter dependent state $\hat{\rho}(\theta)=\hat{U}(\theta)\hat{\rho}\hat{U}(\theta)^{\dagger}$. For unitary evolutions, the QFI is independent of $\theta$ and only depends on the initial state $\hat{\rho}$ and the generator $\hat{G}$, which we express by writing $F_{\mathrm{Q}}[\hat{\rho}(\theta)]=F_{\mathrm{Q}}[\hat{\rho},\hat{G}$]. In this case, Eq. (28) reduces to
$F_{\mathrm{Q}}[\hat{\rho},\hat{G}]=2\sum_{k,l}\frac{(\lambda_{k}-\lambda_{l})^{2}}{\lambda_{k}+\lambda_{l}}|\langle k|\hat{G}|l\rangle|^{2}\;.$ (30)
@@
Gaussian states. In addition, an analytical expression for the QFI can also be obtained for arbitrary Gaussian states. For a single-mode Gaussian state $\hat{\rho}(\theta)$ under unitary evolution, the QFI is given in terms of their first-moment vector $\langle\mathbf{\hat{r}}\rangle_{\hat{\rho}(\theta)}$ and covariance matrix $\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}$] as [33, 75, 76, 58]
$F_{Q}[\hat{\rho}(\theta)]=\frac{1}{2}\frac{\,\mathrm{Tr}\left[(\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\frac{\partial}{\partial\theta}\Gamma[\hat{\rho}(\theta),\mathbf{\hat{r}}])^{2}\right]}{1+\mathcal{P}(\hat{\rho}(\theta))^{2}}+\left(\frac{\partial\langle\mathbf{\hat{r}}\rangle_{\hat{\rho}(\theta)}}{\partial\theta}\right)^{T}\Gamma^{-1}[\hat{\rho}(\theta),\mathbf{\hat{r}}]\left(\frac{\partial\langle\mathbf{\hat{r}}\rangle_{\hat{\rho}(\theta)}}{\partial\theta}\right),$ (33)
@@
is the purity. For an evolution generated by a single-mode quadratic Hamiltonian of the form (19), we obtain
$F_{Q}[\hat{\rho},\frac{1}{2}\hat{\bm{r}}^{T}H\hat{\bm{r}}$] $=\frac{\,\mathrm{Tr}\left[\Gamma[\hat{\rho},\mathbf{\hat{r}}]H\right]^{2}-4\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]\det{H}}{\frac{1}{4}+\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}+\frac{\langle\hat{\bm{r}}\rangle_{\hat{\rho}}^{T}H\Gamma[\hat{\rho},\mathbf{\hat{r}}]H\langle\hat{\bm{r}}\rangle_{\hat{\rho}}}{\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}$. (35)
A derivation of this result is given in App. F. For mixed states under non-unitary Gaussian channels (II.3), the purity-changing term $2\left(\frac{\partial\mathcal{P}(\hat{\rho}(\theta))}{\partial\theta}\right)^{2}/(1-\mathcal{P}(\hat{\rho}(\theta))^{4})$ is added to the QFI (33[33, 58].
Convexity. The QFI is a convex function of the quantum state, i.e. for any probability distribution $p_{k}$ and family of quantum states $\hat{\rho}_{k}$, we have
@@ III.2 Method of moments
Figure 3: Illustration of the method of moments. First, the calibration curve $f(\theta)$ relating $\langle\hat{M}\rangle_{\hat{\rho}(\theta)}$ to $\theta$ is recorded or predicted from theory. Then, from $\mu$ measurements of the observable $\hat{M}$ the sample average $\overline{M}_{\mu}$ is computed and subsequently used to estimate the unknown parameter by inverting the calibration curve, i.e. $\theta_{\text{est}}=f^{-1}(\overline{M}_{\mu})$. In the central limit, when $\mu\gg 1$, the distribution of the sample average becomes Gaussian with variance $(\Delta\hat{M})^{2}_{\hat{\rho}(\theta)}/\mu$. The error on the estimator is then determined by maximum likelihood estimation based on the sample average data, or equivalently, by error propagation, and depends on the derivative of the calibration curve.
The QFI describes the maximal metrological sensitivity that can be extracted from the quantum state $\hat{\rho}$ using an optimal estimation strategy. A practical approach to parameter estimation based on the average value of a single observable $\hat{M}$ is given by the method of moments, see Fig. 3. This method provides a simple estimator for $\theta$ by comparing the sample average of $\mu$ repeated measurements of $\hat{M}$ with a previously established relation $\langle\hat{M}\rangle_{\hat{\rho}(\theta)}$, i.e. a “calibration curve”. Once this relation is known, implementation of the method of moments is straightforward: The estimator is simply given by the parameter value that yields equality between the sample average and the calibration curve.
@@
This maximum is achieved by the symmetric logarithmic derivative $\hat{L}$, defined in Eq. (29[77].
Dichotomic measurements. For dichotomic measurements the method of moments equals the classical Fisher information. Consider an observable $\hat{M}=m_{1}\hat{\Pi}_{1}+m_{2}\hat{\Pi}_{2}$ with two eigenvalues $m_{i}$ and corresponding projectors $\hat{\Pi}_{i}$, such that $\hat{\Pi}_{1}+\hat{\Pi}_{2}=\hat{\mathbb{I}}$. The conditional probability is given by $P(m_{1}|\theta)=\,\mathrm{Tr}\left[\hat{\rho}(\theta)\hat{\Pi}_{1}\right]=1-\,\mathrm{Tr}\left[\hat{\rho}(\theta)\hat{\Pi}_{2}\right]=1-P(m_{2}|\theta)$. The variance of $\hat{M}$ reads
@@
$F[\hat{\rho}(\theta),\{\hat{E}_{m}\}]$ $=\frac{1}{P(m_{1}|\theta)}\left(\frac{\partial P(m_{1}|\theta)}{\partial\theta}\right)^{2}+\frac{1}{P(m_{2}|\theta)}\left(\frac{\partial P(m_{2}|\theta)}{\partial\theta}\right)^{2}$